# Li Zhou1, Lutong Yu1, Dongchu Xie1, Shaohuan Cheng3, Wenyan Li4, Haizhou Li1, 2

### v2

@@ Abstract
Culture is a rich and dynamic domain that evolves across both geography and time. However, existing studies on cultural understanding with vision-language models (VLMs) primarily emphasize geographic diversity, often overlooking the critical temporal dimensions. To bridge this gap, we introduce Hanfu-Bench, a novel, expert-curated multimodal dataset. Hanfu, a traditional garment spanning ancient Chinese dynasties, serves as a representative cultural heritage that reflects the profound temporal aspects of Chinese culture while remaining highly popular in Chinese contemporary society. Hanfu-Bench comprises two core tasks: cultural visual understanding and cultural image transcreation. The former task examines temporal-cultural feature recognition based on single- or multi-image inputs through multiple-choice visual question answering, while the latter focuses on transforming traditional attire into modern designs through cultural element inheritance and modern context adaptation. Our evaluation shows that closed VLMs perform comparably to non-experts on visual cutural understanding but fall short by 10% to human experts, while open VLMs lags further behind non-experts. For the transcreation task, multi-faceted human evaluation indicates that the best-performing model achieves a success rate of only 42%. Our benchmark provides an essential testbed, revealing significant challenges in this new direction of temporal cultural understanding and creative adaptation.  [^1] Following Jacovi et al. (2023), the Hanfu-Bench dataset is publicly available at lizhou21/Hanfu-Bench under the CC BY-NC-SA 4.0 License. The code details are freely available for reuse at lizhou21/TemporalCulture.
@@
Li Zhou1thanks: Corresponding author, Lutong Yu1, Dongchu Xie1, Shaohuan Cheng3, Wenyan Li4, Haizhou Li1, 2 1The Chinese University of Hong Kong, Shenzhen, 2Shenzhen Research Institute of Big Data, 3 Chengdu Technological University, 4University of Copenhagen lizhou21@cuhk.edu.cn
**footnotetext: Equal contribution.
@@ 1 Introduction
[…]
@@ 3.3 Expert Annotation
[…]
@@ 4.1 VQA Tasks and Formualtion
We developed a rule-based pipeline to automatically generate questions for both SVQA and MVQA tasks. For all questions, answer candidates are selected to ensure (1) exactly one correct answer exists and (2) no duplicate options appear. For SVQA, we create question templates across eight feature categories from our annotations, with answer candidates drawn from the annotated features. For MVQA, we designed templates to generate questions about features across image groups. For example, as shown in Figure 2, when asking "Which image belongs to a period after Tang Dynasty?", we filter images based on their “period" annotations to meet our answer selection criteria. The complete set of base questions and the corresponding attribute values are provided in Appendix A.5.
@@ Acknowledgments
This work is supported by Shenzhen Science and Technology Program (Shenzhen Key Laboratory, Grant No. ZDSYS20230626091302006), Shenzhen Science and Technology Research Fund (Fundamental Research Key Project, Grant No.JCYJ20220818103001002), and Program for Guangdong Introducing Innovative and Entrepreneurial Teams, Grant No. 2023ZT10X044. Wenyan Li is supported by the Lundbeck Foundation (BrainDrugs grant: R279-2018-1145).
@@ 2. Annotation Details:
    Progress Tracking: The "Progress" metric represents the ratio of completed pages to the total number of datasets. The "Current Progress" indicates the sequence of the current image within the dataset. To proceed to the next image, it is mandatory to complete all questions on the current page. The use of "Jump" is discouraged to prevent data transmission failure.
@@
    Image Preservation Decision: Based on the image quality and content, determine whether the image can reflect the characteristics of the Hanfu style. If the image quality is low or the style features are extremely unclear, select "Not Retained".
@@
    Other Labels: Refer to the "Various Hanfu Style Features" in the appendix for judgment. Note that if the Hanfu in the image is not a " lianchang", then " lianchang sleeve", " lianchang collar", and " lianchang lapel" should all be selected as "No". If uncertain, choose "Unable to Judge".
@@
    Specific Hanfu Elements Identification: For the question "What specific Hanfu elements does this piece of clothing have?", consider the clothes appearing on the current page. Despite potential slight differences in style and color, focus on several more obvious and distinctive Hanfu elements.
@@
    Data Preservation: After completing all questions on the page, it is imperative to click "Save Labels" before proceeding to the next page to avoid data transmission issues.
@@ A.5 Base question templates and feature value
Type {CJK}UTF8gbsn传统汉服形制 Traditional Hanfu style, 汉服改良版 Improved Han clothing, 汉元素服饰 Han-element clothing
Gender {CJK}UTF8gbsn 男 Male, 女 Female
Period {CJK}UTF8gbsn秦汉时期 Qin and Han Dynasties, 魏晋时期 Wei and Jin Dynasties, 唐朝 Tang Dynasty, 宋朝 Song Dynasty. 明朝 Ming Dynasty
Sleeve {CJK}UTF8gbsn窄袖 Narrow sleeves, 直袖 Straight sleeves, 半袖 Half sleeves, 琵琶袖 Pipa sleeves, 垂胡袖 Chuihu sleeves, 大袖 Large sleeves
Jin {CJK}UTF8gbsn大襟 Da Jin, 对襟 Dui Jin, 绕襟 Rao Jin
Collar {CJK}UTF8gbsn直领 Straight collar, 坦领 Tan collar, 圆领 Round collar, 方领 Square collar, 立领 Standing collar, 交领 Cross collar
Bottoms {CJK}UTF8gbsn 破群Po qun, 裤 Pants, 马面裙 Mamian qun, 褶裙 Zhe qun
Outerwear {CJK}UTF8gbsn 比甲 Bijia, 半臂 Banbi, 云肩 Yunjian, 褙子 Beizi, 披帛 Pibo, 披风 Pifeng
@@
Type {CJK}UTF8gbsn图片中的服饰通常属于以下哪个类型?Which type does the clothing in this image typically belong to, among the following options?
Gender {CJK}UTF8gbsn图片中的服饰通常适合什么性别?Which gender is the clothing in this image typically suitable for, among the following options?
Period {CJK}UTF8gbsn图片中的服饰属于以下哪个时期的风格?Which historical period does the style of the clothing in this image belong to, among the following options?
Sleeve {CJK}UTF8gbsn图片中服饰的袖子属于以下哪种类型?Which type does the sleeves of the clothing in this image belong to, among the following options?
Jin {CJK}UTF8gbsn图片中服饰的襟型属于以下哪种类型?Which type does the lapel style of the clothing in this image belong to, among the following options?
Collar {CJK}UTF8gbsn图片中服饰的领型属于以下哪种类型?Which type does the collar style of the clothing in this image belong to, among the following options?
Bottoms {CJK}UTF8gbsn图片中服饰的下身是什么类型的?Which type does the lower garment of the clothing in this image belong to, among the following options?
Outerwear {CJK}UTF8gbsn图片中服饰的外搭是什么?Which type does the outer layer of the clothing in this image belong to, among the following options?
@@
Type {CJK}UTF8gbsn以下图片中的服饰属于汉服改良版的是?Which of the following pictures of clothing belongs to the Improved Han clothing?
{CJK}UTF8gbsn以下图片中的服饰属于传统汉服形制的是?Which of the following images represents a traditional Hanfu style?
{CJK}UTF8gbsn以下图片中不属于汉服改良版的服饰是?Which of the following pictures of clothing does not belong to the improved Han clothing?
@@
{CJK}UTF8gbsn以下图片中不属于传统汉服形制的服饰是? Which of the following images does NOT belong to traditional Hanfu styles?
{CJK}UTF8gbsn以下图片中的服饰属于汉元素服饰的是?Which of the following images represents Han element clothing?
@@
{CJK}UTF8gbsn以下图片中形制与其他服饰不同的是? Which of the following images shows a different structural style compared to the other garments?
@@
{CJK}UTF8gbsn以下图片中风格与其他服饰不同的有?Which of the following images shows a different overall style compared to the others?
Gender {CJK}UTF8gbsn以下图片中的服饰属于男性服饰的是? Which of the following pictures of clothing belongs to male’s clothing?
{CJK}UTF8gbsn以下图片中不属于男性服饰的是?Which of the following images does NOT depict male attire?
{CJK}UTF8gbsn以下图片中的服饰属于女性服饰的是?Which of the following images depicts female’s Hanfu attire?
{CJK}UTF8gbsn以下图片中不属于女性服饰的是?Which of the following images does NOT show women’s attire?
@@
{CJK}UTF8gbsn以下图片中服饰性别属性与其他图片不同的是?Which of the following images shows different gender attributes compared to the others?
@@
Period {CJK}UTF8gbsn以下图片中属于唐朝以后时期的服饰有? Which of the following images shows clothing styles from after the Tang Dynasty?
@@
{CJK}UTF8gbsn以下图片中属于魏晋时期以后时期的服饰有? Which of the following images shows clothing styles from after the Wei-Jin period?
@@
{CJK}UTF8gbsn以下的服装中比以上图片中的服饰更古老的有?Which of the following garments is older than those shown in the images above?
{CJK}UTF8gbsn以下图片中的服饰属于秦汉时期的风格的有?Which of the following images represents Qin and Han Dynasties’ style clothing?
@@
{CJK}UTF8gbsn以下图片中的服饰属于明朝的风格的有?Which of the following images represents Ming Dynasty-style clothing?
@@
{CJK}UTF8gbsn以下图片中不属于唐朝的服饰有?Which of the following images does NOT show Tang Dynasty attire?
@@
{CJK}UTF8gbsn以下图片中不属于宋朝的服饰有?Which of the following images does NOT represent Song Dynasty-era clothing?
@@
{CJK}UTF8gbsn以下图片中不属于明朝的服饰有?Which of the following images does NOT show Ming Dynasty-era attire?
{CJK}UTF8gbsn以下图片中的服饰属于魏晋时期的风格的有? Which of the following images represents Wei-Jin period style clothing?
@@
{CJK}UTF8gbsn以下图片中的服饰属于唐朝的风格的有? Which of the following images represents Tang Dynasty-style clothing?
@@
{CJK}UTF8gbsn以下图片中属于宋朝以后时期的服饰有? Which of the following images displays clothing styles from after the Song Dynasty?
@@
{CJK}UTF8gbsn以下图片中的服饰属于宋朝的风格的有? Which of the following images represents Song Dynasty-style clothing?
@@
Sleeve {CJK}UTF8gbsn以下图片中服饰袖型与其他图片不同的是? Which of the following images has a sleeve style that differs from the others?
{CJK}UTF8gbsn以下图片中服饰袖型属于垂胡袖的是? Which of the following images features the ’Chuihu sleeve’?
{CJK}UTF8gbsn以下图片中服饰袖型不属于琵琶袖的是?Which of the following images does NOT feature pipa sleeves?
{CJK}UTF8gbsn以下图片中服饰袖型属于直袖的是? Which of the following images displays straight sleeves?
{CJK}UTF8gbsn以下图片中服饰袖型属于窄袖的是? Which of the following images features narrow sleeves?
{CJK}UTF8gbsn以下图片中服饰袖型不属于直袖的是? Which of the following images does NOT have straight sleeves?
{CJK}UTF8gbsn以下图片中服饰袖型属于琵琶袖的是? Which of the following images displays the ’Pipa sleeve’?
{CJK}UTF8gbsn以下图片中服饰袖型不属于窄袖的是? Which of the following images does NOT have narrow sleeves?
{CJK}UTF8gbsn以下图片中服饰袖型不属于大袖的是? Which of the following images does NOT feature large sleeves?
{CJK}UTF8gbsn以下图片中服饰袖型属于半袖的是? Which of the following images features half sleeves?
{CJK}UTF8gbsn以下图片中服饰袖型属于大袖的是? Which of the following images shows large sleeves?
@@
Collar {CJK}UTF8gbsn以下图片中服饰领型属于交领的是? Which of the following images shows the ’Cross collar’ neckline style?
@@
{CJK}UTF8gbsn以下图片中服饰领型不属于直领的是? Which of the following images does NOT display the ’straight collar’ neckline?
{CJK}UTF8gbsn以下图片中服饰领型属于直领的是? Which of the following images shows a straight collar neckline?
@@
{CJK}UTF8gbsn以下图片中服饰领型与其他图片不同的是? Which of the following images has a neckline style that differs from the others?
{CJK}UTF8gbsn以下图片中服饰领型属于立领的是? Which of the following images has a standing collar?
{CJK}UTF8gbsn以下图片中服饰领型属于圆领的是? Which of the following images features a round collar neckline?
{CJK}UTF8gbsn以下图片中服饰领型不属于立领的是? Which of the following images does NOT have a standing collar?
{CJK}UTF8gbsn以下图片中服饰领型属于方领的是? Which of the following images features a square collar neckline?
{CJK}UTF8gbsn以下图片中服饰领型属于坦领的是? Which of the following images has a ’Tan collar’?
{CJK}UTF8gbsn以下图片中服饰领型不属于圆领的是? Which of the following images does NOT have a round collar?
{CJK}UTF8gbsn以下图片中服饰领型不属于交领的是? Which of the following images does NOT have a cross-collar?
@@
Jin {CJK}UTF8gbsn以下图片中服饰襟型与其他图片不同的是? Which of the following images shows a different garment closure type compared to the others?
@@
{CJK}UTF8gbsn以下图片中服饰襟型不属于大襟的是? Which of the following images does NOT feature the ’Dajin’ closure style?
@@
{CJK}UTF8gbsn以下图片中服饰襟型不属于对襟的是? Which of the following images does NOT show the ’Duijin’ closure style?
@@
{CJK}UTF8gbsn以下图片中服饰襟型属于绕襟的是? Which of the following images features a wrapped closure (’Raojin’) style?
@@
{CJK}UTF8gbsn以下图片中服饰襟型属于对襟的是? Which of the following images shows a symmetrical front closure (’Duijin’)?
{CJK}UTF8gbsn以下图片中服饰襟型属于大襟的是? Which of the following images shows the ’Dajin’?
@@
Bottoms {CJK}UTF8gbsn以下图片中下半身服饰种类属于马面裙的是? Among the following pictures which type of lower body clothing belongs to ’Mamian Qun?’
@@
{CJK}UTF8gbsn以下图片中下半身服饰种类属于裤的是? Which of the following images features lower-body clothing categorized as trousers?
@@
{CJK}UTF8gbsn以下图片中下半身服饰种类不属于马面裙的是? Which of the following images does NOT depict a ’Mamian qun’ as the lower-body attire?
@@
{CJK}UTF8gbsn以下图片中下半身服饰种类与其他图片不同的是? Which of the following images shows a different lower-body garment style compared to the others?
@@
{CJK}UTF8gbsn以下图片中下半身服饰种类属于破群的是? Which of the following images depicts a ’Poqun’ as the lower-body garment?
{CJK}UTF8gbsn以下图片中下半身服饰种类属于褶裙的是? Which of the following images depicts pleated skirts (’Zhequn’)?
@@
{CJK}UTF8gbsn以下图片中下半身服饰种类不属于褶裙的是? Which of the following images does NOT depict pleated skirts (’Zhequn’)?
@@
Outerwear {CJK}UTF8gbsn以下图片中外搭服饰种类属于披帛的是? Which of the following types of outerwear in the pictures belongs to the category of ’Pibo’?
@@
{CJK}UTF8gbsn以下图片中外搭服饰种类不属于披帛的是? Which of the following images shows an outer garment that is NOT a ’Pibo’?
@@
{CJK}UTF8gbsn以下图片中外搭服饰种类不属于褙子的是? Which of the following images features an outer garment that is NOT a ’Beizi’?
@@
{CJK}UTF8gbsn以下图片中外搭服饰种类与其他图片不同的是? Which of the following images features an outer garment that differs from the others?
{CJK}UTF8gbsn以下图片中外搭服饰种类不属于云肩的是? Which of the following images does NOT include a ’Yunjian’?
{CJK}UTF8gbsn以下图片中外搭服饰种类不属于披风的是? Which of the following images does NOT include a ’Pifeng’ as outerwear?
{CJK}UTF8gbsn以下图片中外搭服饰种类属于比甲的是? Which of the following images includes a ’Bijia’ as outerwear?
{CJK}UTF8gbsn以下图片中外搭服饰种类属于云肩的是? Which of the following images features a ’Yunjian’?
{CJK}UTF8gbsn以下图片中外搭服饰种类属于披风的是? Which of the following images includes a ’Pifeng’?
{CJK}UTF8gbsn以下图片中外搭服饰种类属于半臂的是? Which of the following images includes a ’Banbi’ as outerwear?
{CJK}UTF8gbsn以下图片中外搭服饰种类属于褙子的是? Which of the following images features a ’Beizi’?
{CJK}UTF8gbsn以下图片中外搭服饰种类不属于比甲的是? Which of the following images does NOT include a ’Bijia’?

### v3

@@ Abstract
Culture is a rich and dynamic domain that evolves across both geography and time. However, existing studies on cultural understanding with vision-language models (VLMs) primarily emphasize geographic diversity, often overlooking the critical temporal dimensions. To bridge this gap, we introduce Hanfu-Bench, a novel, expert-curated multimodal dataset. Hanfu, a traditional garment spanning ancient Chinese dynasties, serves as a representative cultural heritage that reflects the profound temporal aspects of Chinese culture while remaining highly popular in Chinese contemporary society. Hanfu-Bench comprises two core tasks: cultural visual understanding and cultural image transcreation. The former task examines temporal-cultural feature recognition based on single- or multi-image inputs through multiple-choice visual question answering, while the latter focuses on transforming traditional attire into modern designs through cultural element inheritance and modern context adaptation. Our evaluation shows that closed VLMs perform comparably to non-experts on visual cutural understanding but fall short by 10% to human experts, while open VLMs lags further behind non-experts. For the transcreation task, multi-faceted human evaluation indicates that the best-performing model achieves a success rate of only 42%. Our benchmark provides an essential testbed, revealing significant challenges in this new direction of temporal cultural understanding and creative adaptation.  [^1] Following Jacovi et al. (2023), the Hanfu-Bench dataset is publicly available at lizhou21/Hanfu-Bench under the CC BY-NC-SA 4.0 License. The code details are freely available for reuse at hlt-cuhksz/TemporalCulture.
@@
Li Zhou1thanks: Corresponding author, Lutong Yu1, Dongchu Xie1, Shaohuan Cheng3, Wenyan Li4, Haizhou Li1,2 1The Chinese University of Hong Kong, Shenzhen 2Shenzhen Research Institute of Big Data, 3Chengdu Technological University 4University of Copenhagen lizhou21@cuhk.edu.cn, lutongyu@link.cuhk.edu.cn
[…]
@@ 1 Introduction
Traditional Chinese Hanfu exhibits distinct characteristics in collar styles, sleeve designs, garment structures, and layering combinations across different historical periods Yu (2024); Liu et al. (2022); Cai and Liu (2024); Kang et al. (2024), as illustrated in Figure 1.222The top part was hand-drawn by one of our authors, who is familiar with Hanfu. For example, Hanfu of the Qin-Han period predominantly adhered to the Shenyi (深衣) one-piece robe system, distinguished by its Raojin (绕襟) construction—a wrapped-collar design with diagonally overlapping garment panels forming a cross-collar closure  Kang et al. (2024). Tang dynasty fashion featured the Pibo (披帛)—a long, lightweight silk scarf draped asymmetrically over robes—serves as a quintessential outer accessory Liu et al. (2024b), while Ming dynasty clothing prominently showcased the Mamian Qun (马面裙) as the characteristic skirt style Bao and Guo (2025); Lin and Niu (2023). These temporal variations of Hanfu highlight the intrinsic complexity of cultural knowledge, requiring an understanding spanning both historical and aesthetic dimensions. This poses significant challenges for visual-language models (VLMs) Chang et al. (2024); Li et al. (2025a, b) in capturing culture-specific features across different historical periods.
@@ 3.3 Expert Annotation
The global features are: Type (categorized as Traditional Hanfu, Improved Han clothing, or Han-element clothing), [^5] Although sourced from high-quality Hanfu retailers, some designs include modern adaptations, aligning with the objectives of Task 2 in our work. Only outfits within these three categories are retained. Period (identifying the dynasty for Traditional Hanfu), and Gender (indicating the intended or suitable gender). The local features focus on specific components of the outfit, including Sleeve, Collar, Jin (襟, similar to lapels), Bottoms, and Outerwear. For each local feature, annotators are asked to choose from predefined options. They are encouraged to provide text descriptions for unique features not covered by these options and may also select “none” or “unsure” where appropriate. We recruit nine annotators with expertise in Hanfu, comprising four authors of this study and five members of a Hanfu club. To ensure thorough and consistent assessments, each outfit is evaluated by three annotators. Detailed guidelines (Appendix A.2) are provided to enhance annotation reliability, including comprehensive explanations of each feature. The meta-information for a Hanfu outfit consists of feature values where all three annotators reached uniform agreement. We mark the feature value as “unsure” when agreement could not be reached.
@@ 4.1 VQA Tasks and Formualtion
We developed a rule-based pipeline to automatically generate questions for both SVQA and MVQA tasks. For all questions, answer candidates are selected to ensure (1) exactly one correct answer exists and (2) no duplicate options appear. For SVQA, we create question templates across eight feature categories from our annotations, with answer candidates drawn from the annotated features. For MVQA, we designed templates to generate questions about features across image groups. For example, as shown in Figure 2, when asking ”Which image belongs to a period after Tang Dynasty?”, we filter images based on their “period” annotations to meet our answer selection criteria. The complete set of base questions and the corresponding attribute values are provided in Appendix A.5.
@@ Acknowledgments
This work is supported by Shenzhen Science and Technology Program (Shenzhen Key Laboratory, Grant No. ZDSYS20230626091302006), Program for Guangdong Introducing Innovative and Entrepreneurial Teams (Grant No. 2023ZT10X044), Shenzhen Stability Science Program 2023, Shenzhen Key Lab of Multi-Modal Cognitive Computing, Shenzhen Science and Technology Research Fund (Fundamental Research Key Project, Grant No.JCYJ20220818103001002), the International Science and Technology Cooperation Center, Ministry of Science and Technology of China (under grant 2024YFE0203000), the Internal Project of Shenzhen Research Institute of Big Data. Wenyan Li is supported by the Lundbeck Foundation (BrainDrugs grant: R279-2018-1145).
@@ 2. Annotation Details:
    Progress Tracking: The ”Progress” metric represents the ratio of completed pages to the total number of datasets. The ”Current Progress” indicates the sequence of the current image within the dataset. To proceed to the next image, it is mandatory to complete all questions on the current page. The use of ”Jump” is discouraged to prevent data transmission failure.
@@
    Image Preservation Decision: Based on the image quality and content, determine whether the image can reflect the characteristics of the Hanfu style. If the image quality is low or the style features are extremely unclear, select ”Not Retained”.
@@
    Other Labels: Refer to the ”Various Hanfu Style Features” in the appendix for judgment. Note that if the Hanfu in the image is not a ” lianchang”, then ” lianchang sleeve”, ” lianchang collar”, and ” lianchang lapel” should all be selected as ”No”. If uncertain, choose ”Unable to Judge”.
@@
    Specific Hanfu Elements Identification: For the question ”What specific Hanfu elements does this piece of clothing have?”, consider the clothes appearing on the current page. Despite potential slight differences in style and color, focus on several more obvious and distinctive Hanfu elements.
@@
    Data Preservation: After completing all questions on the page, it is imperative to click ”Save Labels” before proceeding to the next page to avoid data transmission issues.
@@ A.5 Base question templates and feature value
Type 传统汉服形制 Traditional Hanfu style, 汉服改良版 Improved Han clothing, 汉元素服饰 Han-element clothing
Gender 男 Male, 女 Female
Period 秦汉时期 Qin and Han Dynasties, 魏晋时期 Wei and Jin Dynasties, 唐朝 Tang Dynasty, 宋朝 Song Dynasty. 明朝 Ming Dynasty
Sleeve 窄袖 Narrow sleeves, 直袖 Straight sleeves, 半袖 Half sleeves, 琵琶袖 Pipa sleeves, 垂胡袖 Chuihu sleeves, 大袖 Large sleeves
Jin 大襟 Da Jin, 对襟 Dui Jin, 绕襟 Rao Jin
Collar 直领 Straight collar, 坦领 Tan collar, 圆领 Round collar, 方领 Square collar, 立领 Standing collar, 交领 Cross collar
Bottoms 破群Po qun, 裤 Pants, 马面裙 Mamian qun, 褶裙 Zhe qun
Outerwear 比甲 Bijia, 半臂 Banbi, 云肩 Yunjian, 褙子 Beizi, 披帛 Pibo, 披风 Pifeng
@@
Type 图片中的服饰通常属于以下哪个类型?Which type does the clothing in this image typically belong to, among the following options?
Gender 图片中的服饰通常适合什么性别?Which gender is the clothing in this image typically suitable for, among the following options?
Period 图片中的服饰属于以下哪个时期的风格?Which historical period does the style of the clothing in this image belong to, among the following options?
Sleeve 图片中服饰的袖子属于以下哪种类型?Which type does the sleeves of the clothing in this image belong to, among the following options?
Jin 图片中服饰的襟型属于以下哪种类型?Which type does the lapel style of the clothing in this image belong to, among the following options?
Collar 图片中服饰的领型属于以下哪种类型?Which type does the collar style of the clothing in this image belong to, among the following options?
Bottoms 图片中服饰的下身是什么类型的?Which type does the lower garment of the clothing in this image belong to, among the following options?
Outerwear 图片中服饰的外搭是什么?Which type does the outer layer of the clothing in this image belong to, among the following options?
@@
Type 以下图片中的服饰属于汉服改良版的是?Which of the following pictures of clothing belongs to the Improved Han clothing?
以下图片中的服饰属于传统汉服形制的是?Which of the following images represents a traditional Hanfu style?
以下图片中不属于汉服改良版的服饰是?Which of the following pictures of clothing does not belong to the improved Han clothing?
@@
以下图片中不属于传统汉服形制的服饰是? Which of the following images does NOT belong to traditional Hanfu styles?
以下图片中的服饰属于汉元素服饰的是?Which of the following images represents Han element clothing?
@@
以下图片中形制与其他服饰不同的是? Which of the following images shows a different structural style compared to the other garments?
@@
以下图片中风格与其他服饰不同的有?Which of the following images shows a different overall style compared to the others?
Gender 以下图片中的服饰属于男性服饰的是? Which of the following pictures of clothing belongs to male’s clothing?
以下图片中不属于男性服饰的是?Which of the following images does NOT depict male attire?
以下图片中的服饰属于女性服饰的是?Which of the following images depicts female’s Hanfu attire?
以下图片中不属于女性服饰的是?Which of the following images does NOT show women’s attire?
@@
以下图片中服饰性别属性与其他图片不同的是?Which of the following images shows different gender attributes compared to the others?
@@
Period 以下图片中属于唐朝以后时期的服饰有? Which of the following images shows clothing styles from after the Tang Dynasty?
@@
以下图片中属于魏晋时期以后时期的服饰有? Which of the following images shows clothing styles from after the Wei-Jin period?
@@
以下的服装中比以上图片中的服饰更古老的有?Which of the following garments is older than those shown in the images above?
以下图片中的服饰属于秦汉时期的风格的有?Which of the following images represents Qin and Han Dynasties’ style clothing?
@@
以下图片中的服饰属于明朝的风格的有?Which of the following images represents Ming Dynasty-style clothing?
@@
以下图片中不属于唐朝的服饰有?Which of the following images does NOT show Tang Dynasty attire?
@@
以下图片中不属于宋朝的服饰有?Which of the following images does NOT represent Song Dynasty-era clothing?
@@
以下图片中不属于明朝的服饰有?Which of the following images does NOT show Ming Dynasty-era attire?
以下图片中的服饰属于魏晋时期的风格的有? Which of the following images represents Wei-Jin period style clothing?
@@
以下图片中的服饰属于唐朝的风格的有? Which of the following images represents Tang Dynasty-style clothing?
@@
以下图片中属于宋朝以后时期的服饰有? Which of the following images displays clothing styles from after the Song Dynasty?
@@
以下图片中的服饰属于宋朝的风格的有? Which of the following images represents Song Dynasty-style clothing?
@@
Sleeve 以下图片中服饰袖型与其他图片不同的是? Which of the following images has a sleeve style that differs from the others?
以下图片中服饰袖型属于垂胡袖的是? Which of the following images features the ’Chuihu sleeve’?
以下图片中服饰袖型不属于琵琶袖的是?Which of the following images does NOT feature pipa sleeves?
以下图片中服饰袖型属于直袖的是? Which of the following images displays straight sleeves?
以下图片中服饰袖型属于窄袖的是? Which of the following images features narrow sleeves?
以下图片中服饰袖型不属于直袖的是? Which of the following images does NOT have straight sleeves?
以下图片中服饰袖型属于琵琶袖的是? Which of the following images displays the ’Pipa sleeve’?
以下图片中服饰袖型不属于窄袖的是? Which of the following images does NOT have narrow sleeves?
以下图片中服饰袖型不属于大袖的是? Which of the following images does NOT feature large sleeves?
以下图片中服饰袖型属于半袖的是? Which of the following images features half sleeves?
以下图片中服饰袖型属于大袖的是? Which of the following images shows large sleeves?
@@
Collar 以下图片中服饰领型属于交领的是? Which of the following images shows the ’Cross collar’ neckline style?
@@
以下图片中服饰领型不属于直领的是? Which of the following images does NOT display the ’straight collar’ neckline?
以下图片中服饰领型属于直领的是? Which of the following images shows a straight collar neckline?
@@
以下图片中服饰领型与其他图片不同的是? Which of the following images has a neckline style that differs from the others?
以下图片中服饰领型属于立领的是? Which of the following images has a standing collar?
以下图片中服饰领型属于圆领的是? Which of the following images features a round collar neckline?
以下图片中服饰领型不属于立领的是? Which of the following images does NOT have a standing collar?
以下图片中服饰领型属于方领的是? Which of the following images features a square collar neckline?
以下图片中服饰领型属于坦领的是? Which of the following images has a ’Tan collar’?
以下图片中服饰领型不属于圆领的是? Which of the following images does NOT have a round collar?
以下图片中服饰领型不属于交领的是? Which of the following images does NOT have a cross-collar?
@@
Jin 以下图片中服饰襟型与其他图片不同的是? Which of the following images shows a different garment closure type compared to the others?
@@
以下图片中服饰襟型不属于大襟的是? Which of the following images does NOT feature the ’Dajin’ closure style?
@@
以下图片中服饰襟型不属于对襟的是? Which of the following images does NOT show the ’Duijin’ closure style?
@@
以下图片中服饰襟型属于绕襟的是? Which of the following images features a wrapped closure (’Raojin’) style?
@@
以下图片中服饰襟型属于对襟的是? Which of the following images shows a symmetrical front closure (’Duijin’)?
以下图片中服饰襟型属于大襟的是? Which of the following images shows the ’Dajin’?
@@
Bottoms 以下图片中下半身服饰种类属于马面裙的是? Among the following pictures which type of lower body clothing belongs to ’Mamian Qun?’
@@
以下图片中下半身服饰种类属于裤的是? Which of the following images features lower-body clothing categorized as trousers?
@@
以下图片中下半身服饰种类不属于马面裙的是? Which of the following images does NOT depict a ’Mamian qun’ as the lower-body attire?
@@
以下图片中下半身服饰种类与其他图片不同的是? Which of the following images shows a different lower-body garment style compared to the others?
@@
以下图片中下半身服饰种类属于破群的是? Which of the following images depicts a ’Poqun’ as the lower-body garment?
以下图片中下半身服饰种类属于褶裙的是? Which of the following images depicts pleated skirts (’Zhequn’)?
@@
以下图片中下半身服饰种类不属于褶裙的是? Which of the following images does NOT depict pleated skirts (’Zhequn’)?
@@
Outerwear 以下图片中外搭服饰种类属于披帛的是? Which of the following types of outerwear in the pictures belongs to the category of ’Pibo’?
@@
以下图片中外搭服饰种类不属于披帛的是? Which of the following images shows an outer garment that is NOT a ’Pibo’?
@@
以下图片中外搭服饰种类不属于褙子的是? Which of the following images features an outer garment that is NOT a ’Beizi’?
@@
以下图片中外搭服饰种类与其他图片不同的是? Which of the following images features an outer garment that differs from the others?
以下图片中外搭服饰种类不属于云肩的是? Which of the following images does NOT include a ’Yunjian’?
以下图片中外搭服饰种类不属于披风的是? Which of the following images does NOT include a ’Pifeng’ as outerwear?
以下图片中外搭服饰种类属于比甲的是? Which of the following images includes a ’Bijia’ as outerwear?
以下图片中外搭服饰种类属于云肩的是? Which of the following images features a ’Yunjian’?
以下图片中外搭服饰种类属于披风的是? Which of the following images includes a ’Pifeng’?
以下图片中外搭服饰种类属于半臂的是? Which of the following images includes a ’Banbi’ as outerwear?
以下图片中外搭服饰种类属于褙子的是? Which of the following images features a ’Beizi’?
以下图片中外搭服饰种类不属于比甲的是? Which of the following images does NOT include a ’Bijia’?